A closed tank contains 100 lbs of air at a pressure of 200 psia and a temperaure of 180 0F. The air is subsequently cooled to 80 0F. Find the: a.) the initial specific volume, ft3/lbm; b.) the volume of the tank, ft3; c.) the final specific volume, ft3/lbm; d.) the final pressure, psia.
1
Expert's answer
2020-05-15T09:47:22-0400

Pressure= 200 psia, Air present =100 lbs,"T_1= 180 ^oF, T_2= 80^oF,"

(a) Specific Volume,


"P_1V_1=RT_1"


"200\times V= 53.33\times (180+460)"


"V=170.656" "ft^3"/lbm

(b)

Volume of the tank ,


V= "\frac {170.656 \times 100}{60}=284.426 ft^3"

(c) for the final specific volume


"P_2V_2=RT_2"


"V_2= \frac{ 53..33\times (80+460)}{200}=143.991"

(d)

Final pressure


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


P2=333.33 psia